OpenStudy (anonymous):

Find the intercepts and relative extrema for the graph of the function. 1. f(x)=1/4(x+2)(x-1)^2

OpenStudy (anonymous):

for x-intercept solve: 1/4(x+2)(x-1)^2=0 for y- intercept solve: 1/4(0+2)(0-1)^2=f(0) for extrema find derivative f'(x) and solve f'(x)=0

OpenStudy (anonymous):

for x-inter I got -2 ,1 for y-inter I got 1/2 but im confused how to find the extrema

OpenStudy (anonymous):

1º find the derivative

OpenStudy (anonymous):

use the rule of derivative of product: (uv)'=u'v+uv' where u=1/4(x+2) and v=(x-1)^2

OpenStudy (anonymous):

((1/4)(x+2)(x-1)^2)=(1/4)(x+2)(x-1)^2+(1/4)(x+2)(x-1)^2

OpenStudy (anonymous):

((1/4)(x+2)(x-1)^2)=1/2 (-1+x)^2 (2+x) ??

OpenStudy (anonymous):

u'=1/4 v'=2(x-1)=2x-2 So: (uv)'=u'v+uv'=1/4(x-1)^2+1/4(x+2)2(x-1)=1/4(x-1)(x-1+2x+4)=1/4(x-1)(3x+3) now set it to 0: 1/4(x-1)(3x+3)=0 so at x=1 and x=-1 are the extrema. To find the actual value of the extrema plug this x values nto original equation and find f(x)